If there are 0.045 mols of NH3 in 23.6 mL of solution, what is its concentration? Also,
what is the pH?
10.)
1
Expert's answer
2020-05-03T15:25:04-0400

Molar concentration = n(NH3)/V(solution) = 0.045mol/0.0236L = 1.907 mol/L

NH3 + H2O = NH4+ + OH-

C(OH-) = 1.907 M

pOH = - lg[OH-]

pH = 14 - pOH